OpenStudy (anonymous):

Simplify: x^(1/3) times x^(3/5) Please show solving process. Thanks!

OpenStudy (anonymous):

1/3 + 3/5 = 5/15 + 9/15 = 14/15 so x^14/15 I think

OpenStudy (anonymous):

well as u can see x is the same so x^(1/3+3/5) so x^(5+3/15) x^8/15